OpenStudy (anonymous):

2x+3 x -----=--- 4 4 Can someone explain the steps for this

OpenStudy (shamil98):

Cross multiply, what do you get?

OpenStudy (shamil98):

You could also multiply the value of (4) to both sides to get rid of the 4 in the denominators in both sides leaving the equation to just be 2x +3 = x

OpenStudy (texaschic101):

This is a proportion, so you cross multiply (2x + 3)/4 = x/4 (4)(x) = (4)(2x + 3) 4x = 8x + 12 4x - 8x = 12 -4x = 12 x = -3 check... (2x + 3)/4 = x/4 (2(-3) + 3)/4 = -3/4 (-6 + 3)/4 = -3/4 -3/4 = -3/4 (correct) x = -3

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

do you have any questions ?

jhonyy9 (jhonyy9):

@texaschic101 so when you have two fractions with the same denominators what mean this ?

OpenStudy (shamil98):

If they are equal to each other like \[\huge \frac{ 2x+3 }{ 4 } = \frac{ x }{ 4 }\] Then the numerators are equal to each other You can eliminate the denominators because: \[\huge 4 \times \frac{ 2x +3 }{ 4 } = \frac{ x }{ 4 } \times 4\] \[\huge 2x + 3 = x\]

OpenStudy (shamil98):

I should say if they are equal to each other and have the same denominator*

jhonyy9 (jhonyy9):

nice @shamil98 but do it till the end

jhonyy9 (jhonyy9):

so than x= ?

jhonyy9 (jhonyy9):

x=-3 or not ?

OpenStudy (texaschic101):

I did it a little different......but the answer will be the same. I crossed multiplied, and then distributed through the parenthesis, then solved for x.

OpenStudy (shamil98):

yeah x is still -3 i just reduced it.